### v1

# Interpretation of a Discrete de Rham method as a Finite Element System

Abstract
We show that the DDR method can be interpreted as defining a computable consistent discrete $\mathrm{L}^{2}$ product on a conforming FES defined by PDEs. Without modifying the numerical method itself, this point of view provides an alternative approach to the analysis. The conformity and consistency properties we prove are stronger than those previously shown. We can also recover some of the other results that have been proved about DDR, from those that have already been proved, in principle, in the general context of FES. We also bring VEM, the Virtual Element Method, into the discussion.
MSC: 65N30, 58A12.
Keywords: finite elements, differential forms.
Introduction
We give a new presentation of the DDR method of [27, 9]. We interpret it as a numerical method based on discrete spaces defined piecewise by solving certain new Partial Differential Equations (PDEs), on each cell of a mesh. As such, the method is consistent in the sense of Strang’s lemmas, and conforming in the sense of producing subcomplexes of natural $\mathrm{L}^{2}$-based Hilbert complexes. Even though the solutions to these new local PDEs are not computable, the numerical method itself is ”fully discrete” (computable). This type of consistency and conformity are stronger than those previously proved for DDR. The interpretation also embeds the method in the Finite Element System framework, and allows to leverage the results already obtained in that setting.
The framework of Finite Element Systems (FES) was introduced in [11] and expanded upon in [20, §5] (this independent section is available on arXiv [^1] S. H. Christiansen, Finite element systems of differential forms, arXiv:1006.4779, 2015.) and subsequent papers. It is an environment in which to define and analyze sequences of spaces of differential forms, that are finite elements in a broad sense. The underlying mesh can consist of cells (in particular polytopes) and the basis functions can be general (not necessarily based on polynomials). Thus one may call them polytopal discrete de Rham sequences, though this expression is given an a priori different meaning below. These spaces were conforming in the sense that the fields were $\mathrm{L}^{2}$-integrable with exterior derivative also $\mathrm{L}^{2}$-integrable, which enforces that traces, in the sence of pullbacks to faces, coincide (they are ”single-valued”). This encodes a type of partial continuity that is common for the mixed finite elements of electromagnetics and fluid flow [8], such as those of Raviart-Thomas-Nédélec. The latter were extended to differential forms in [32]. A synthesis is provided in Finite Element Exterior Calculus (FEEC) [3].
Since then the framework of FES has been generalized to accomodate higher regularity of the fields [18] (as in so-called ”Stokes complexes” relevant to fluid flow) and other types of fields [19] (e.g. symmetric tensors and their associated curvature tensors, as in Riemannian geometry). The definition has a natural expression in the language of categories and a short summary is provided in [15]. It now covers all finite elements described according to Ciarlet’s definition, including non-conforming ones (such as Morley and Crouzeix-Raviart elements).
For the purposes of this paper, however, the definition given in [11] is sufficient. The goal back then was to define a discrete de Rham finite element sequence on polytopal meshes that mimicked the lowest order mixed finite elements that correspond to Whitney forms. Indeed these were known only for simplices (and products of simplices handled by tensor product constructions). For this presentation will also refer to the more comprehensive presentation in [20, §5]. Short presentations can also be found in later publications, in particular [21, §2].
A high order version of the construction of [11] yielding minimal spaces (sometimes referred to as ”serendipity spaces”) was given in [12, 16]. The method of [11, 16] was conceptually based on harmonic extensions of differential forms, which is not in general computable. But with the help of, for instance, mesh refinements and known discrete spaces on the refinement, adequate discrete harmonic extensions were defined, yielding ”fully discrete” numerical methods with algebraic and analytical properties identical with those of standard mixed finite elements.
By modifying the harmonic extension PDE that underlied the above constructions, one can obtain upwinded complexes, to treat convection dominated diffusion problems [13, 17]. The framework has also been used to give compact descriptions of several known finite element methods [16, 21]. The generalized Whitney forms introduced in [11] have also been used to analyze Discrete Exterior Calculus (DEC) in [30].
The Virtual Element Method (VEM), reviewed in [5], has been developed with similar goals as FES and some similar basic principles, such as definition by recursive extension of the fields, from low dimensional cells. But, crucially, no submesh is used in VEM to define the discretization spaces or the numerical method. The challenge is to reconstruct enough information on the fields from the degrees of freedom, to obtain stable and consistent approximations of the bilinear forms defining the partial differential equation to solve. A connection between the mixed VEM (defined in dimension 3) of [4] and FES (in arbitrary dimension) was made in [16, §2.1] and will be used later.
The Discrete de Rham (DDR) method was introduced in dimension 2 and 3 in [27] and has since been further developed in several papers. We refer to [9] for an overview and a description of the method valid for differential forms of arbitrary degree in arbitrary dimension.
The DDR construction was inspired at least in part by Hybrid High Order (HHO) methods. Accordingly, it has been presented as a non-conforming method involving spaces of non-matching fields on cells of all dimension. Non-conformity of the spaces poses major challenges for the analysis of the method, many of which have been addressed in the literature [25, 26]. In particular an original notion of ”adjoint consistency” has been developed.
The close relationship between DDR and VEM has already been pointed out in [6]. In some versions these methods can be seen to have the same degrees of freedom, and they coincide with those of FES provided in [16, §2.1].
In this paper, we reinterpret DDR as a conforming FES equipped with computable consistent approximations of the $\mathrm{L}^{2}$ scalar products. The degrees of freedom of [16, §2.1] are used. This shows that these examples of DDR, VEM and FES are very similar, if not identical, in practice. They seem to differ mainly in terms of interpretation and analysis. The main novelty here is just that a new harmonic extension PDE, modifying the one used in [16, §2.1], is introduced to define the discrete spaces.
This point of view allows us to recover some of the results that have been proved for VEM or DDR [25, 26], as special cases of results already proved, in their principle, in the general context of FES. To sum up:
  • Consistency of the discrete $\mathrm{L}^{2}$ scalar product now follows from the fact that the new local PDE allows the explicit recursive computation of $\mathrm{L}^{2}$ products with enough polynomials, from the degrees of freedom.
  • Cohomological properties developed for FES, especially [20, Proposition 5.15], now carry over to DDR. This passage from local exactness (on each cell of the mesh), to globally correct cohomology (on the computational domain) illustrates that the existence of unisolvent degrees of freedom on a FES can be interpreted as a softness condition on sheaves (see [33, §III.2] for definitions).
  • Uniformly bounded commuting projections, building on [34, 10], have been described for FEEC [3, 23] and FES [20, §5.3]. They enable one to recover Poincaré-Friedrich estimates, which are basic to the stability analysis of mixed methods.
In the setting of FES one furthermore obtains refinements:
  • Translation estimates are given in [20, Proposition 5.71], and have applications to the equations of fluid flow.
  • Discrete Rellich compactness, as introduced by Kikuchi for Maxwell [7, §IV.19], holds generally, as proved first in [10, Corollary 5] with applications to the Hodge-Laplace eigenvalue problem on manifolds. See also [24] on this topic.
  • Discrete Sobolev injections hold [20, Propositions 5.69, 5.70] and have applications to, for instance, non-linear Maxwell wave equations as in [22]. See also [31] on this topic.
  • Fractional order Sobolev estimates are provided in [14], with applications to the Dirac equation.
Remark 0.1.
More recent constructions of local bounded commuting projections [29, 2] carry over to FES in a very natural way. The method of [29] extends to FES of differential forms and the method of [2] extends to general FES as in [19] (the latter method does not appeal to a wedge product). In fact the FES framework seems to identify exactly the algebraic properties of discretization spaces needed for these constructions. A general analysis of stability and compactness properties of these methods remains, in the context of FES, to be fleshed out. See [28] for the state of the art.
While we have little to add to the design of the DDR and VEM methods (a part from a new rationale for neglecting certain terms), we hope the FES point of view can contribute usefully to their analysis. We also emphasize that many of the technical computations that appear here without further ado, already appear exemplified in the above papers on DDR and VEM, though perhaps not in this order and generality.
1 Notations
Let $\mathcal{T}$ be a polytopal mesh of some Euclidean polyhedral domain. When considering a polytope $T$ with a face $T^{\prime}$, if $u$ is a differential form on $T$, by its trace on $T^{\prime}$ we always mean the trace in the sense of pullback of $u$ to $T^{\prime}$ by the inclusion of $T^{\prime}$ in $T$.
For each $k\in\mathbb{N}$ and $T\in\mathcal{T}$ we let $X^{k}(T)$ be the space of differential $k$-forms on $T$ which are in $\mathrm{L}^{2}(T)$ with exterior derivative in $\mathrm{L}^{2}(T)$. The $\mathrm{L}^{2}$ norm is denoted $|\cdot|$ and the natural graph norm on $X^{k}(T)$ is denoted $\|\cdot\|$.
The subspace of $X^{k}(T)$ consisting of forms with 0 traces on $\partial T$ is denoted $X^{k}_{0}(T)$. Subscript 0 will have the same meaning for other spaces as well. Elements on $X^{k}(T)$ have well defined traces, in negative index Sobolev spaces, only on codimension one faces $T^{\prime}\subseteq\partial T$. However, if these traces in turn are imposed to be in $X^{k}(T^{\prime})$, one can repeat the procedure and follow traces all the way down. One gets a space of forms with improved boundary regularity and some nice properties, such as [21, Proposition 2.4]. All our discrete spaces will have this type of extra boundary regularity.
The space of Whitney $k$-forms (which could also be attributed to Weil and de Rham) was originally presented in terms of partitions of unity, such as barycentric coordinates, but it can also be presented in terms of the Koszul operator [3, §3.2]. For each cell $T\in\mathcal{T}$, we let $b_{T}$ be its barycentre (we suppose it is ”computable”). On $T$, the Koszul operator $\kappa_{T}$ is the contraction by the vector field on $T$ given as $x\mapsto x-b_{T}$ (for $x\in T$). We adopt notations from [3]. The original lowest order Whitney forms are then $\mathcal{P}_{\!-}^{1}\Lambda^{k}(T)$. The high order version are denoted $\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$:
$\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)=\{u\in\mathcal{P}^{r}\Lambda^{k}(T)\ :\ \kappa_{T}u\in\mathcal{P}^{r}\Lambda^{k-1}(T)\}.$ (1)
These spaces were defined in [32] in terms of the standard Poincaré null-homotopy operators rather than the Koszul operators and were described in FES terms in [21] under the name ”trimmed” polynomial differential forms.
The formal adjoint of $\mathrm{d}$ with respect to the $\mathrm{L}^{2}$ scalar product is denoted $\mathrm{d}^{\star}$. The Hodge star operator is denoted $\scriptstyle\bigstar$}. We have $\mathrm{d}^{\star}=\pm{\scriptstyle\bigstar}\mathrm{d}{\scriptstyle\bigstar}$.
A third point of view on Whitney forms, developed in [11], is that they are the solutions to the PDE system:
$\mathrm{d}^{\star}\mathrm{d}u=0,\quad\mathrm{d}^{\star}u=0,$ (2)
supposed to hold on simplices of all dimensions (equipped with translation invariant metrics). For 0-forms this reduces to $\Delta u=0$. For $k$-forms on simplices (or flat cells for that matter) of dimension $k$, this reduces to $\mathrm{d}^{\star}u=0$, which is a characterization of the constant $k$-forms.
Even though we adopt the FES point of view, we try to express our results first without reference to those definitions. On the other hand [21, §2] contains a glossary and a polished short presentation.
2 Lowest order case
Let $T$ be a cell. The case $\dim T=0$ need not be detailed, so we suppose that $\dim T=d>0$.
In this section we define $W^{k}(T)$ as a slight modification of $\mathcal{P}_{\!-}^{1}\Lambda^{k}(T)$, namely, that for, $k=0$ we impose that the elements of $W^{0}(T)$ should have integral 0.
Then we define:
$Z^{k}_{0}(T)$ $=\{u\in X^{k}_{0}(T)\ :\ \forall w\in W^{d-k}(T)\quad\int_{T}u\wedge w=0\}.$ (3)
Proposition 2.1.
The spaces $Z^{\scriptscriptstyle{\bullet}}_{0}(T)$ form a sequence which is exact except at index $k=\dim T$, where the integral determines an isomorphism from the cohomology group to $\mathbb{R}$.
Proof.
We have a short exact sequence of complexes:
$0\to Z^{\scriptscriptstyle{\bullet}}_{0}(T)\to X^{\scriptscriptstyle{\bullet}}_{0}(T)\to W^{d-{\scriptscriptstyle{\bullet}}}(T)^{\star}\to 0.$ (4)
The sequence $W^{\scriptscriptstyle{\bullet}}(T)$ is exact, including at indices 0 and $d$. Hence $Z^{\scriptscriptstyle{\bullet}}_{0}(T)$ inherits the cohomology of $X^{\scriptscriptstyle{\bullet}}_{0}(T)$, by the snake lemma, in its full version [1, Theorem 3.5 page 600]. ∎
We define:
$Y^{k}(T)=\{u\in Z^{k}_{0}(T)\ :\ u\perp\mathrm{d}Z^{k-1}_{0}(T)\}.$ (5)
Orthogonality is taken with respect to the $\mathrm{L}^{2}$ scalar product (also elsewhere in this paper). We remark that Proposition 2.1 entails that the following is an isomorphism:
$\mathrm{d}:Y^{k-1}(T)\to\ker\mathrm{d}|_{Z^{k}_{0}(T)},$ (6)
for $k<d$, and that, for the case $k=d$, we also have the isomorphism:
$\mathrm{d}:Y^{d-1}(T)\to\ker\textstyle\int_{T}(\cdot)|_{Z^{d}(T)}.$ (7)
We define $B^{k}(T)$ to be the subspace of $X^{k}(T)$ consisting of fields $u\in X^{k}(T)$ such that the following system of PDEs holds:
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a+\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=0,$ (8)
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0,\ \int_{T}u\wedge\kappa_{T}b^{\prime}$ $=0,$ (9)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=0,$ (10)
for all $u^{\prime}\in X^{k}_{0}(T)$, for some $a\in\mathrm{P}^{0}\Lambda^{d-k}(T)$, some $b\in\mathrm{P}^{0}\Lambda^{d-k+1}(T)$, and some $v\in Y^{k-1}(T)$, for all $v^{\prime}\in Y^{k-1}(T)$, all $a^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k}(T)$ and all $b^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k+1}(T)$.
Notice that $a^{\prime},b^{\prime}$ and $v^{\prime}$ enforce constraints on $u$, and that $a,b,v$ can be interpreted as the corresponding Lagrange multipliers. We refer to $a^{\prime}$ and $b^{\prime}$ as the finite dimensional constraints.
Remark 2.1.
Removing the finite dimensional space of constraints represented by $a,a^{\prime}$ and $b,b^{\prime}$, and allowing general $v,v^{\prime}$, we obtain the harmonic extension used in [11]. Indeed the equations considered there were:
$\mathrm{d}^{\star}\mathrm{d}u+\mathrm{d}v$ $=0,$ (11)
$\mathrm{d}^{\star}u$ $=0.$ (12)
There, discrete variants obtained from weak formulations on suitable finite dimensional spaces, were also used to make the method ”fully discrete”. The Lagrange multiplier $v$ was always 0.
We now detail the special cases $k=0$ and $k=d$ of the definition of $B^{k}(T)$.
For $k=0$ we are left with the following system:
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a=0,$ (13)
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0.$ (14)
For $k=d$, we are left with:
$\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v=0,$ (15)
$\int_{T}u\wedge\kappa_{T}b^{\prime}=0,$ (16)
$\int_{T}u\cdot\mathrm{d}v^{\prime}=0,$ (17)
Notice that $\kappa_{T}b$ is a general element of $W^{0}(T)$. It has integral 0 because of the choice of origin in $\kappa_{T}$. Let $\overline{u}$ be the average of $u$. Then $u-\overline{u}$ is in $Y^{d}(T)$ and has integral 0, so it is 0. That is $u=\overline{u}$. In other words $B^{d}(T)$ consists of the constant $d$-forms.
Then we define $A^{k}(T)$ to be the subspace of $X^{k}(T)$ consisting of fields $u$ whose trace on any face $T^{\prime}$ of $T$ (including $T^{\prime}=T$) is in $B^{k}(T^{\prime})$.
We make the following claims, which we will justify subsequently:
  • A field in $B^{k}(T)$, for $k<\dim T$ is uniquely determined by its boundary traces. Moreover the extension problem defined by the above PDE is wellposed.
  • The family of spaces $A^{k}(T)$ is closed under traces and exterior derivative, making them a FES.
  • The spaces $A^{k}(T)$ are finite dimensional and have one degree of freedom per $k$-face of $T$. This ensures softness of the FES.
  • For each $T\in\mathcal{T}$, the spaces $A^{\scriptscriptstyle{\bullet}}(T)$ form an exact sequence resolving $\mathbb{R}$. Together with softness this ensures compatibility of the FES.
  • The space $A^{k}(T)$ contains the Whitney forms $\mathcal{P}_{\!-}^{1}\Lambda^{k}(T)$. This ensures approximation properties.
  • For any $u\in A^{k}(T)$ and any $a\in\mathrm{P}^{0}\Lambda^{d-k}(T)$, the integral $\int_{T}u\wedge a$ can be computed exactly from the degrees of freedom, in a recursive way.
According to plan, we first comment on the well-posedness of the PDE. We will use the following.
Proposition 2.2.
On a cell $T$ of dimension $d$ and for $k<d$, the above system of PDEs for $B^{k}(T)$ is wellposed, for given boundary data.
Proof.
Extend the boundary data to an arbitrary field $\tilde{u}\in X^{k}(T)$ (for instance by ordinary harmonic extension). Then write the unknown $u\in X^{k}(T)$ as $u=\tilde{u}+\hat{u}$ with $\hat{u}\in X^{k}_{0}(T)$. We now consider the system:
$\int_{T}\mathrm{d}\hat{u}\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a+\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=-\int_{T}\mathrm{d}\tilde{u}\cdot\mathrm{d}u^{\prime},$ (18)
$\int_{T}\mathrm{d}\hat{u}\wedge\kappa_{T}a^{\prime}=-\int_{T}\mathrm{d}\tilde{u}\wedge\kappa_{T}a^{\prime},\ \int_{T}\hat{u}\wedge\kappa_{T}b^{\prime}$ $=-\int_{T}\tilde{u}\wedge\kappa_{T}b^{\prime},$ (19)
$\int_{T}\hat{u}\cdot\mathrm{d}v^{\prime}$ $=-\int_{T}\tilde{u}\cdot\mathrm{d}v^{\prime},$ (20)
Since $k<d$ we can now integrate by parts the constraints, as in:
$\int_{T}\mathrm{d}\hat{u}\wedge\kappa_{T}a^{\prime}+\int_{T}\hat{u}\wedge\kappa_{T}b^{\prime}$ $=\int_{T}\hat{u}\wedge((-1)^{k+1}\mathrm{d}\kappa_{T}a^{\prime}+\kappa_{T}b^{\prime}),$ (21)
$=\int_{T}\hat{u}\wedge(\frac{(-1)^{k+1}}{d-k}a^{\prime}+\kappa_{T}b^{\prime}),$ (22)
For $k=0$ there is no $b^{\prime}$ term. We recognize the general expression of a $w^{\prime}\in\mathrm{W}^{d-k}(T)$:
$w^{\prime}=\frac{(-1)^{k+1}}{d-k}a^{\prime}+\kappa_{T}b^{\prime},$ (23)
uniquely corresponding to $a^{\prime}$ and $b^{\prime}$.
We can rewrite the system for $u\in X^{k}_{0}(T)$ as, for all $u^{\prime}\in X^{k}_{0}(T)$:
$\int_{T}\mathrm{d}\hat{u}\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=-\int_{T}\mathrm{d}\tilde{u}\cdot\mathrm{d}u^{\prime},$ (24)
$\int_{T}\hat{u}\wedge w^{\prime}$ $=-\int_{T}\mathrm{d}\tilde{u}\wedge\kappa_{T}a^{\prime}-\int_{T}\tilde{u}\wedge\kappa_{T}b^{\prime},$ (25)
$\int_{T}\hat{u}\cdot\mathrm{d}v^{\prime}$ $=-\int_{T}\tilde{u}\cdot\mathrm{d}v^{\prime},$ (26)
for some $w\in W^{d-k}(T)$ and $v\in Y^{k-1}(T)$, and for all $w^{\prime}\in W^{d-k}(T)$ (written as (23) ) and $v^{\prime}\in Y^{k-1}(T)$.
We look first at the homogeneous problem (zero right sides). The second equation says that $\hat{u}\in Z^{k}_{0}(T)$ and the third one then says that $\hat{u}\in Y^{k}(T)$. The first equation with $u^{\prime}=\hat{u}$ gives $\mathrm{d}\hat{u}=0$. This shows that $\hat{u}=0$. The homogeneous problem thus has only the trivial solution.
The proof of the well-posedness in the PDE sense, is contained in the Lemmas 2.4 and 2.5, below. ∎
We begin with the following remark.
Lemma 2.3.
We have: $({\scriptstyle\bigstar}W^{d-k}(T))\cap\mathrm{d}Y^{k-1}(T)=0$.
Proof.
It holds, more generally, that for $u\in Z^{k}_{0}(T)$ we have $u\perp{\scriptstyle\bigstar}W^{d-k}(T)$ by definition. ∎
Lemma 2.4 (Inf-Sup condition).
Fix $k$ and $T$. Suppose that $E$ is a finite dimensional subspace of $X^{k}(T)$ and that $E\cap\mathrm{d}Y^{k-1}(T)=0$. Then:
$\inf_{v\in\mathrm{d}Y^{k-1}(T),w\in E}\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v\rangle+\langle u,w\rangle|}{\|u\|(\|v\|+\|w\|)}>0.$ (27)
Proof.
If not, construct sequences $v_{n}\in\mathrm{d}Y^{k-1}_{0}(T)$ and $w_{n}\in E$ such that $\|v_{n}\|+\|w_{n}\|=1$ and:
$\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v_{n}\rangle+\langle u,w_{n}\rangle|}{\|u\|}\to 0.$ (28)
Extracting subsequences we suppose futhermore that $v_{n}\to v$ weakly in $X^{k}_{0}(T)$ and that $w_{n}\to w$ strongly in $E$ (by finite dimensionality of $E$).
From the above we get that $v+w=0$. Also $\mathrm{d}Y^{k-1}(T)$ is closed in $X^{k}_{0}(T)$, so $v\in\mathrm{d}Y^{k-1}(T)$. Therefore $v=0$ and $w=0$. Thus we conclude that $w_{n}\to 0$ strongly. From that we get:
$\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v_{n}\rangle|}{\|u\|}\to 0.$ (29)
But this is false, since we can use $u=v_{n}$ in this $\sup$. ∎
Lemma 2.5 (Coercivity on the kernel).
There is $C>0$ such that for all $u\in Z^{k}_{0}(T)$, such that $u\perp\mathrm{d}Y^{k-1}(T)$ (w.r.t. $\mathrm{L}^{2}$ scalar product):
$|u|\leq C|\mathrm{d}u|.$ (30)
Proof.
If not choose a sequence $u_{n}\in Z^{k}_{0}(T)$, such that $u_{n}\perp\mathrm{d}Y^{k-1}(T)$, $|u_{n}|=1$ and $|\mathrm{d}u_{n}|\to 0$. Extracting subsequences we suppose that $(u_{n})$ converges weakly in $\mathrm{L}^{2}$ to some $u$. Then $u\in Z^{k}_{0}(T)$ and $\mathrm{d}u=0$ and $u\perp\mathrm{d}Y^{k-1}(T)$. Hence $u=0$.
We now show that the convergence $u_{n}\to 0$ is actually strong in $\mathrm{L}^{2}$. For that purpose we control $\mathrm{d}^{\star}u_{n}$ in $\mathrm{L}^{2}$. We decompose, with respect to the $\mathrm{L}^{2}$ product:
$X_{0}^{k-1}(T)=Z_{0}^{k-1}(T)\oplus F.$ (31)
Any $v\in X_{0}^{k-1}(T)$ such that $v\perp\mathrm{d}X^{k-2}_{0}(T)$, can now accordingly be written $v=v^{\prime}+w$ with $v^{\prime}\in Z^{k-1}_{0}(T)$ and $w\in F$. We have:
$\sup_{v\in X_{0}^{k-1}(T)}\frac{|\int_{T}u_{n}\cdot\mathrm{d}v|}{|v|}\leq\sup_{w\in F}\frac{|\int_{T}u_{n}\cdot\mathrm{d}w|}{|w|}.$ (32)
Since $F$ is finite dimensional and $u_{n}\to 0$ weakly in $\mathrm{L}^{2}$, we get that $\mathrm{d}^{\star}u_{n}\to 0$ strongly in $\mathrm{L}^{2}$.
Since the subspace of $X^{k}_{0}(T)$ of consisting of fields $u$ with $\mathrm{d}^{\star}u\in\mathrm{L}^{2}(T)$, with its natural topology, is compactly injected in $\mathrm{L}^{2}(T)$, $(u_{n})$ converges strongly in $\mathrm{L}^{2}$ norm to 0. But this contradicts $|u_{n}|=1$. ∎
Proposition 2.6 (FES).
The family of spaces $A^{k}(T)$ is closed under traces and the exterior derivative.
Proof.
Closedness under traces follows from the definition of $A^{k}(T)$, that already took into account all the traces.
Closedness under the exterior derivative will follow from the fact that if $u\in B^{k}(T)$ then $\mathrm{d}u\in B^{k+1}(T)$. Suppose then that $u\in B^{k}(T)$. We have for all $u^{\prime}\in X^{k}_{0}(T)$:
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v=0.$ (33)
If now $u^{\prime}\in Y^{k}(T)$, the second and third terms are 0. This gives the last constraint on $\mathrm{d}u$.
The two finite dimensional constraints are trivially satisfied. The top equation is also trivially satisfied, with zero Lagrange multipliers. ∎
Proposition 2.7 (Degrees of Freedom).
On $A^{k}(T)$ the integrals on $k$-faces of $T$ constitute unisolvent degrees of freedom.
Proof.
Any given $u\in A^{k}(T)$ is determined by its codimension one traces, that can be followed all the way down to faces of dimension $k$, where $u$ must be constant, hence determined by its integral. Conversely from any values of these integrals, a $u\in A^{k}(T)$ can be reconstructed by recursive extension of the corresponding constant forms on $k$-cells. ∎
Proposition 2.8 (Local sequence exactness).
The sequence $A^{\scriptscriptstyle{\bullet}}(T)$ is exact (more precisely it resolves $\mathbb{R}$).
Proof.
The degrees of freedom provide a cochain isomorphism with the cellular cochain complex, which is known to be exact on any given cell. ∎
Proposition 2.9 (Approximation).
The spaces $A^{k}(T)$ contain $\mathcal{P}_{\!-}^{1}\Lambda^{k}(T)$.
Proof.
Whitney forms satisfy $\mathrm{d}^{\star}u=0$ and $\mathrm{d}^{\star}\mathrm{d}u=0$, so the crucial points are the finite dimensional constraints.
It was proved in [10, Preprint, Proposition 3.9] that $\wedge:\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)\times\mathcal{P}_{\!-}^{s}\Lambda^{l}(T)\to\mathcal{P}_{\!-}^{r+s}\Lambda^{k+l}(T)$. A simpler proof based on the Koszul operator was then provided in [3, Equation (3.16)]. Based on this result, if $u\in\mathcal{P}_{\!-}^{1}\Lambda^{k}(T)$ and $b^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k+1}(T)$ we have that $u\wedge\kappa_{T}b$ is a linear $d$-form. Its integral can therefore be obtained from evaluation at the barycentre of $T$. Since furthermore $\kappa_{T}$ was constructed with respect to this point, we get that $\int_{T}u\wedge\kappa_{T}b^{\prime}=0$. Similarly $\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0$ for $a^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k}(T)$. ∎
Proposition 2.10 (Computation of moments).
For $u\in A^{k}(T)$, the integrals $\int_{T}u\wedge a^{\prime}$ against constant $a^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k}(T)$ can be computed recursively from the degrees of freedom.
Proof.
Let $T$ be a cell of dimension $d$. The case $k=d$ is immediate, so we suppose $k<d$.
We integrate by parts the first constraint (as before, but now with a boundary term):
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=\int_{\partial T}u\wedge\kappa_{T}a^{\prime}+\int_{T}u\wedge\frac{(-1)^{k+1}}{d-k}a^{\prime}.$ (34)
The left side is zero by definition. On the right side, the second term is of the form that we would like to compute.
We are left with the problem of evaluating the first term on the right side. It will be decomposed into integrals on the (d1)(d-1) dimensional faces $T^{\prime}$ of $T$. That is, we require terms of the form:
$\int_{T^{\prime}}u\wedge\kappa_{T}a^{\prime}.$ (35)
We use that $(\kappa_{T}-\kappa_{T^{\prime}})a^{\prime}$ is constant on $T^{\prime}$ and that by definition:
$\int_{T^{\prime}}u\wedge\kappa_{T^{\prime}}a^{\prime}=0.$ (36)
The boundary term has been reduced to integrals of $u$ wedged against constants.
This method can be pursued, decreasing the dimension by one at each step, all the way down to the case where the integrals against constants are actually degrees of freedom. ∎
For a given $u\in A^{k}(T)$ one can then compute the $\mathrm{L}^{2}$ projection of $u$ onto constants $\mathcal{P}^{0}\Lambda^{k}(T)$, from its degrees of freedom. We denote it as $\gamma_{T}u$. In practice one computes $\gamma_{T^{\prime}}u$ starting with $\dim T^{\prime}=k$ and increasing dimension of $T^{\prime}$ by one at each step.
Following VEM practice one introduces a symmetric bilinear form $s_{h}$ on $A^{k}(T)$, defined from the degrees of freedom, which is bounded and coercive in $\mathrm{L}^{2}$, uniformly with respect to $T$. It should scale like the $\mathrm{L}^{2}$ inner product (under scaling maps $x\to h_{T}x$, where $h_{T}$ is the diameter of $T$) but no consistency is required. One then defines the discrete $\mathrm{L}^{2}$-product on $A^{k}(T)$ by:
$\langle u,u^{\prime}\rangle_{h}=\int_{T}\gamma_{T}u\cdot\gamma_{T}u^{\prime}+s_{h}(u-\gamma_{T}u,u^{\prime}-\gamma_{T}u^{\prime}).$ (37)
The second term on the right side is referred to as the stablizing term.
The crucial points are that this discrete inner product is computable, coercive and consistent. It is consistent in the sense that if at least one of $u,u^{\prime}$ is constant, it gives the exact value:
$\langle u,u^{\prime}\rangle_{h}=\int_{T}u\cdot u^{\prime}.$ (38)
Remark also that the exterior derivative $\mathrm{d}:A^{k}(T)\to A^{k+1}(T)$ is computable from the degrees of freedom on $A^{k}(T)$ to those of $A^{k+1}(T)$. From there, various numerical methods for various PDEs can be defined and analyzed as in VEM.
3 High order case
We fix a polynomial degree $r$ with $r\geq 1$. We will define a method that properly reconstructs $k$-forms that are polynomials of degree $r$. More precisely we will define spaces $A^{k}(T)$ that contain $\mathcal{P}^{r}\Lambda^{k}(T)$ and such that the spaces $\mathcal{P}_{\!-}^{r}\Lambda^{d^{\prime}-k}(T^{\prime})$ (with $d^{\prime}=\dim T^{\prime}$) constitute unisolvent degrees of freedom. The crucial point, in practice, is that we can compute the more general moments against $\mathcal{P}^{r}\Lambda^{d^{\prime}-k}(T^{\prime})$ explicitely (just removing the minus).
Remark 3.1 (Link with low order case).
To a large extent the low order case treated previously would correspond to $r=0$. However the theory is different on some points, because the spaces $\mathcal{P}_{\!-}^{0}\Lambda^{k}(T)$ (which look trivial but are the degrees of freedom of the Whitney forms) lack some of the properties of $\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$, for $r\geq 1$.
We set:
$U^{k}(T)=\{u\in\mathcal{P}_{\!-}^{r+1}\Lambda^{k}(T)\ :\ u\perp\mathcal{P}^{r}\Lambda^{k}(T)\}.$ (39)
Orthogonality is with respect to the $\mathrm{L}^{2}$ product. For instance:
$U^{0}(T)=\{u\in\mathcal{P}^{r+1}(T)\ :\ u\perp\mathcal{P}^{r}(T)\}.$ (40)
In particular this space does not contain any constant function. We also notice that for $d=\dim T$:
$U^{d}(T)=0.$ (41)
Then we change the previous definition we have of $W^{k}(T)$ to the following:
$W^{k}(T)=\mathrm{d}U^{k-1}(T)+U^{k}(T).$ (42)
Proposition 3.1.
We have:
  • The decomposition (42) is $\mathrm{L}^{2}$-orthogonal.
  • The following map is bijective:
    $\left\{\begin{array}[]{rcl}\displaystyle U^{k-1}\times U^{k}&\to&W^{k}(T),\\ \displaystyle(u^{\prime},u)&\mapsto&\mathrm{d}u^{\prime}+u.\end{array}\right$. (43)
  • The sequence $W^{\scriptscriptstyle{\bullet}}(T)$ is exact, including at indices 0 and $d$.
Proof.
(i) Let $u^{\prime}\in U^{k-1}(T)$ then $\mathrm{d}u^{\prime}\in\mathcal{P}^{r}\Lambda^{k}(T)$ so $\mathrm{d}u^{\prime}\perp U^{k}(T)$.
(ii) We also notice that if $\mathrm{d}u^{\prime}=0$ then $u^{\prime}\in\mathcal{P}^{r}\Lambda^{k-1}(T)$ so $u^{\prime}=0$.
(iii) Put now $w=\mathrm{d}u^{\prime}+u$, with $u^{\prime}\in U^{k-1}(T)$ and $u\in U^{k}(T)$. If $\mathrm{d}w=0$ then $\mathrm{d}u=0$. Then $u\in\mathcal{P}^{r}\Lambda^{k}(T)$, so $u=0$. So $w=\mathrm{d}u^{\prime}\in\mathrm{d}W^{k-1}(T)$. ∎
Proposition 3.2.
We have direct sum decompositions:
$\mathcal{P}_{\!-}^{r+1}\Lambda^{k}(T)=\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)\oplus W^{k}(T).$ (44)
Proof.
(i) We make a preliminary observation.
Referring to [3, Theorem 3.1], we write, for $u\in\mathcal{P}^{r+1}\Lambda^{k}(T)$:
[…]
Suppose now that $u\in U^{k}$ and that $\mathrm{d}u\in\mathcal{P}_{\!-}^{r}\Lambda^{k+1}(T)$. We have:
$(k+r+1)u=\mathrm{d}\kappa_{T}u+\kappa_{T}\mathrm{d}u-Q_{r}u.$ (46)
The three terms on the right side are in $\mathcal{P}^{r}\Lambda^{k}(T)$. Hence $u=0$.
(ii) We now show that the sum is direct.
Let $u^{\prime}\in U^{k-1}(T)$ and $u\in U^{k}(T)$ and suppose $w=\mathrm{d}u^{\prime}+u\in\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$. Then $\mathrm{d}u\in\mathcal{P}^{r-1}\Lambda^{k+1}(T)$ so $\mathrm{d}u=0$, so $u=0$. Then $\mathrm{d}u^{\prime}\in\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$ so $u^{\prime}=0$.
(iii) We now show that $\mathcal{P}_{\!-}^{r+1}\Lambda^{k}(T)\subseteq\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)\oplus W^{k}(T)$.
Suppose that $v\in\mathcal{P}_{\!-}^{r+1}\Lambda^{k}(T)$. Write first:
$v=u+v^{\prime},\textrm{with }u\in U^{k}\textrm{ and }v^{\prime}\in\mathcal{P}^{r}\Lambda^{k}(T).$ (47)
Next, as a variant of the above decompositon, write:
[…]
and furthermore:
$\kappa_{T}v^{\prime}=u^{\prime}+v^{\prime\prime},\textrm{with }u^{\prime}\in U^{k-1}\textrm{ and }v^{\prime\prime}\in\mathcal{P}^{r}\Lambda^{k-1}(T).$ (49)
We now have:
$(k+r)v^{\prime}=\mathrm{d}u^{\prime}+\mathrm{d}v^{\prime\prime}+\kappa_{T}\mathrm{d}v^{\prime}-Q_{r-1}v^{\prime}.$ (50)
The first term on the right side is in $\mathrm{d}U^{k-1}$ and the three others are in $\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$. Inserting this in (47) we are done. ∎
We keep the previous definition of $Z^{k}_{0}(T)$, given the new definition of $W^{d-k}(T)$, viz.:
$Z^{k}_{0}(T)=\{u\in X^{k}_{0}(T)\ :\ \forall w\in W^{d-k}(T)\quad\int_{T}u\wedge w=0\}.$ (51)
We remark that $Z^{\scriptscriptstyle{\bullet}}_{0}(T)$ has the same exactness property as in Proposition 2.1, since $W^{\scriptscriptstyle{\bullet}}(T)$ is exact. We similarly define $Y^{k}(T)$ given the new definition of $Z^{k}_{0}(T)$, viz.:
$Y^{k}(T)=\{u\in Z^{k}_{0}(T)\ :\ u\perp\mathrm{d}Z^{k-1}_{0}(T)\}.$ (52)
We define $G^{k}(T)$ as follows:
$G^{k}(T)=\{(f,g)\in\mathcal{P}^{r-1}\Lambda^{k}(T)\times\mathcal{P}^{r-1}\Lambda^{k-1}(T)\ :\ \mathrm{d}^{\star}f=0,\ \mathrm{d}^{\star}g=0\}.$ (53)
A motivating observation for this definition is that for $u\in\mathcal{P}^{r}\Lambda^{k}(T)$ we have:
$(\mathrm{d}^{\star}\mathrm{d}u,\mathrm{d}^{\star}u)\in G^{k}(T).$ (54)
Concerning $k=d$ we note that $\mathrm{d}^{\star}f=0$ means that $f\in\mathcal{P}^{0}\Lambda^{d}(T)\approx\mathbb{R}$.
We define $B^{k}(T)$ to be the subspace of $X^{k}(T)$ consisting of fields $u\in X^{k}(T)$ such that the following system of PDEs holds, for some $(f,g)\in G^{k}(T)$:
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge a+\int_{T}u^{\prime}\wedge b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (55)
$\int_{T}\mathrm{d}u\wedge a^{\prime}=0,\ \int_{T}u\wedge b^{\prime}$ $=0,$ (56)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime},$ (57)
for all $u^{\prime}\in X^{k}_{0}(T)$, for some $a\in U^{d-k-1}(T)$, some $b\in U^{d-k}(T)$, and some $v\in Y^{k-1}(T)$, for all $v^{\prime}\in Y^{k-1}(T)$, all $a^{\prime}\in U^{d-k-1}(T)$ and all $b^{\prime}\in U^{d-k}(T)$.
Remark 3.2 (Old FES interpretation of VEM and DDR).
Removing the finite dimensional space of constraints represented by $a,a^{\prime}$ and $b,b^{\prime}$, and allowing general $v,v^{\prime}$, we obtain the harmonic extension used in [16]. Indeed the equations considered there were:
$\mathrm{d}^{\star}\mathrm{d}u+\mathrm{d}v$ $=f,$ (58)
$\mathrm{d}^{\star}u$ $=g.$ (59)
for some $(f,g)\in G^{k}(T)$. When $k=d$ the top row is replaced by $\int_{T}u=\int_{T}f$.
There, we proved that the spaces of solutions to these PDEs define a FES which is compatible in the sense of [20, Definition 5.12] and that the spaces $\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$ give unisolvent degrees of freedom.
From this point of view, the problem was to show that the discrete $\mathrm{L}^{2}$ product, as defined in [27, 9], is consistent with the true $\mathrm{L}^{2}$ product, on this FES, in the sense of Strang’s lemmas on variational crimes. We could not settle this question. This paper provides a modification of the PDEs leading to an alternative approach: the PDE is more elaborate but consistency is easy to prove, due to Proposition 3.10 below.
For the case $k=0$ of the definition of $B^{k}(T)$ we have that the system becomes:
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge a$ $=\int_{T}f\cdot u^{\prime},$ (60)
$\int_{T}\mathrm{d}u\wedge a^{\prime}$ $=0.$ (61)
For some $a\in U^{1}(T)$ and all $a^{\prime}\in U^{1}(T)$.
Concerning the case $k=d$ we have:
Proposition 3.3.
We have:
$B^{d}(T)=\mathcal{P}^{r}\Lambda^{d}(T).$ (62)
Proof.
The system becomes:
$\int_{T}u^{\prime}\wedge b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (63)
$\int_{T}u\wedge b^{\prime}$ $=0,$ (64)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime},$ (65)
for all $u^{\prime}\in X^{d}(T)$, some $b\in U^{0}(T)$ and all $b^{\prime}\in U^{0}(T)$ , some $y\in Y^{d-1}(T)$ and all $y^{\prime}\in Y^{d}(T)$.
Let $\overline{u}$ be the $\mathrm{L}^{2}$ projection of $u$ on $\mathcal{P}^{r}\Lambda^{d}(T)$. Due to the second equation (64) we have:
$\int_{T}(u-\overline{u})\wedge b^{\prime}=0,$ (66)
for all $b^{\prime}\in\mathcal{P}^{r+1}(T)$. We can choose $t\in X^{d-1}_{0}(T)$ such that $\mathrm{d}t=u-\overline{u}$ and for all $c^{\prime}\in\mathcal{P}_{\!-}^{r+1}\Lambda^{1}(T)$:
$\int_{T}t\wedge c^{\prime}=0.$ (67)
In particular $t\in Z^{d-1}_{0}(T)$.
Since $\mathrm{d}^{\star}g=0$, the last equation (65) is actually true for any $v^{\prime}\in Z^{d-1}_{0}(T)$. Using it for $v^{\prime}=t$ we get:
$\int_{T}|\mathrm{d}t|^{2}=\int_{T}g\cdot t=\pm\int_{T}t\wedge{\scriptstyle\bigstar}g=0.$ (68)
We get $u=\overline{u}$, so $u\in\mathcal{P}^{r}\Lambda^{d}(T)$.
This is also sufficient for the existence of a corresponding $(f,g)\in G^{d}(T)$, since one can choose $f=0$ and $g=\mathrm{d}^{\star}u$, and remark that $u$ solves the system with $b=0$ and $v=0$. ∎
Proposition 3.4.
For $k=d$ the system (55) can be solved iff $f=0$. In that case $u$ is determined up to a constant.
Proof.
The first equation can be rewritten:
$\pm{\scriptstyle\bigstar}b+\mathrm{d}v=f.$ (69)
Integrating this we get $\int_{T}f=0$. Since $f$ is constant (by the condition $\mathrm{d}^{\star}f=0$), we get $f=0$.
On the other hand, if $f=0$, we can find $u\in\mathcal{P}^{r}\Lambda^{d}(T)$ such that $\mathrm{d}^{\star}u=g$. Then $u$ solves the system with $b=0$ and $v=0$.
We notice that the constants solve the homogeneous problem. Suppose now that $u$ solves the homogeneous problem and $\int_{T}u=0$. Since $u\in Z^{d}(T)$ from the second equation, we have $u\in\mathrm{d}Z^{d-1}_{0}(T)=\mathrm{d}Y^{d-1}(T)$. Using $v^{\prime}\in Y^{d-1}(T)$ such that $u=\mathrm{d}v^{\prime}$ in the last equation gives $u=0$. ∎
Proposition 3.5.
For $k<d$ the above PDE system is wellposed and has a unique solution $u$, given boundary data for $u$ and any right sides $f\in\mathrm{L}^{2}\Lambda^{k}(T)$ and $g\in\mathrm{L}^{2}\Lambda^{k-1}(T)$ with $\mathrm{d}^{\star}f=0$ and $\mathrm{d}^{\star}g=0$.
Proof.
As before (by abstract arguments involving compactness) the inf-sup condition and the coercivity on the kernel hold, for this saddlepoint problem. ∎
We define $A^{k}(T)$ as before from $B^{k}(T)$ viz.:
$A^{k}(T)=\{u\in X^{k}(T)\ :\ \forall T^{\prime}\subseteq T\quad u|_{T^{\prime}}\in B^{k}(T^{\prime})\}.$ (70)
It is understood that $T^{\prime}\subseteq T$ means that $T^{\prime}\in\mathcal{T}$ is a face of $T$.
We proceed to show that the spaces $A^{k}(T)$ constitute a FES with some very good properties.
Proposition 3.6 (FES).
The family of spaces $A^{k}(T)$ is closed under traces and the exterior derivative.
Proof.
As before, given $u\in B^{k}(T)$ we prove that $\mathrm{d}u\in B^{k+1}(T)$. Given $u^{\prime}\in Y^{k}(T)$ we have, since the Lagrange multipliers cancel on $Y^{k}(T)$:
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}=\int_{T}f\cdot u^{\prime}.$ (71)
The two finite dimensional constraints required on $\mathrm{d}u$ are also satisfied. ∎
Proposition 3.7 (Approximation).
The space $A^{k}(T)$ contains $\mathcal{P}^{r}\Lambda^{k}(T)$.
Proof.
Given $u\in\mathcal{P}^{r}\Lambda^{k}(T)$ we define $f=\mathrm{d}^{\star}\mathrm{d}u$ and $g=\mathrm{d}^{\star}u$. Then the PDE system is solved with zero Lagrange multipliers. ∎
Proposition 3.8.
For $k<d$ the map $G^{k}(T)\to A^{k}_{0}(T)$ obtained from solving (55) with homogeneous boundary conditions for $u$ is injective.
Proof.
Suppose that the solution to (55) with $u|_{\partial T}=0$ is $u=0$. We then get:
$\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (72)
$0$ $=\int_{T}g\cdot v^{\prime},$ (73)
for all $u^{\prime}\in X^{k}_{0}(T)$, for some $w\in W^{d-k}(T)$, all $w^{\prime}\in W^{d-k}(T)$, some $v\in Y^{k-1}(T)$ and all $v^{\prime}\in Y^{k-1}(T)$.
With $u^{\prime}\in Y^{k}(T)$ we get:
$\int_{T}f\cdot u^{\prime}=0.$ (74)
This also holds for $u^{\prime}=\mathrm{d}v^{\prime}$ for $v^{\prime}\in Y^{k-1}(T)$. We deduce that for all $u^{\prime}\in Z^{k}_{0}(T)$:
$\int_{T}f\cdot u^{\prime}=0.$ (75)
By a density argument with respect to $\mathrm{L}^{2}$ norm, this can be extended to any $u^{\prime}\in\mathrm{L}^{2}\Lambda^{k}(T)$ such that:
$\forall w^{\prime}\in W^{d-k}(T)\quad\int_{T}u^{\prime}\wedge w^{\prime}=0,\textrm{i.e. }u^{\prime}\perp{\scriptstyle\bigstar}W^{d-k}(T).$ (76)
This implies that $f\in{\scriptstyle\bigstar}W^{d-k}(T)$. But in view of (44) we have ${\scriptstyle\bigstar}W^{d-k}(T)\cap\mathcal{P}^{r-1}\Lambda^{k}(T)=0$, so we get $f=0$.
The proof that $g=0$ is identical. ∎
Proposition 3.9 (Degrees of freedom).
Any element $u\in A^{k}(T)$ is uniquely determined by the following degrees of freedom:
$\int_{T^{\prime}}u\wedge e,\quad e\in\mathcal{P}_{\!-}^{r}\Lambda^{d^{\prime}-k}(T^{\prime}),\ d^{\prime}=\dim T^{\prime}.$ (77)
Proof.
The case $k=d$ is immediate. We proceed with the case $k<d$. Recall the notation that $A^{k}_{0}(T)=A^{k}(T)\cap X^{k}_{0}(T)$. We prove that the following bilinear form on $A^{k}_{0}(T)\times\mathcal{P}_{\!-}^{r}\Lambda^{d-k}(T)$ is invertible, in several steps:
$(u,e)\mapsto\int_{T}u\wedge e.$ (78)
(i) For $u\in A^{k}_{0}(T)$ we obtain $u\in Z^{k}_{0}(T)$ from the finite dimensional constraints, and, for all $u^{\prime}\in Z^{k}_{0}(T)$, and all $v^{\prime}\in Y^{k-1}(T)$:
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (79)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime}.$ (80)
We use $u^{\prime}=\mathrm{d}v$ and use $\mathrm{d}^{\star}f=0$ to deduce $\mathrm{d}v=0$ so $v=0$.
(ii) Suppose now that for all $e\in\mathcal{P}_{\!-}^{r}\Lambda^{d-k}(T)$:
$\int_{T}u\wedge e=0.$ (81)
Using (44) we then get, actually, that for all $e\in\mathcal{P}_{\!-}^{r+1}\Lambda^{d-k}(T)$:
$\int_{T}u\wedge e=0.$ (82)
In particular:
$\int_{T}f\cdot u=0.$ (83)
From the choice $u^{\prime}=u$ in the first equation, this gives $\mathrm{d}u=0$. For the case $k=0$ we then deduce $u=0$.
(iii) For $0<k<d$ we proceed as follows.
If, in addition $k>1$, we remark that for any $v^{\prime}\in Z^{k-1}_{0}(T)$ such that $\mathrm{d}v^{\prime}=0$ there is $v^{\prime\prime}\in Z^{k-2}_{0}(T)$ such that $v^{\prime}=\mathrm{d}v^{\prime\prime}$. For such $v^{\prime}$ we have:
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=0,$ (84)
$\int_{T}g\cdot v^{\prime}$ $=\int_{T}g\cdot\mathrm{d}v^{\prime\prime}=0.$ (85)
Combining this with (80) we conclude that for all $v^{\prime}\in Z^{k-1}_{0}(T)$:
$\int_{T}u\cdot\mathrm{d}v^{\prime}=\int_{T}g\cdot v^{\prime}.$ (86)
If $k=1$ there is no $v^{\prime}\in Z^{k-1}_{0}(T)$ such that $\mathrm{d}v^{\prime}=0$, so the above conclusion also holds.
We introduce $\tilde{Z}^{k}_{0}(T)$ defined by:
$\tilde{Z}^{k}_{0}(T)=\{u\in X^{k}_{0}(T)\ :\ \forall e\in\mathcal{P}_{\!-}^{r+1}\Lambda^{d-k}(T)\quad\int_{T}u\wedge e=0\}.$ (87)
We have $\tilde{Z}^{k}_{0}(T)\subseteq Z^{k}_{0}(T)$. The complex $\tilde{Z}^{\scriptscriptstyle{\bullet}}_{0}(T)$ is exact, again by the snake lemma and the exactness properties of $\mathcal{P}_{\!-}^{r+1}\Lambda^{{\scriptscriptstyle{\bullet}}}(T)$, including at index $k=d$.
Since $\mathrm{d}u=0$ and (82) holds we deduce that there is $t\in\tilde{Z}^{k-1}_{0}(T)$ such that $\mathrm{d}t=u$. This entails that for all $e\in\mathcal{P}_{\!-}^{r+1}\Lambda^{d-k+1}(T)$:
$\int_{T}t\wedge e=0.$ (88)
Then we have, setting $u=\mathrm{d}t$ and $v^{\prime}=t$ in (86):
$\int_{T}\mathrm{d}t\cdot\mathrm{d}t=\int_{T}g\cdot t=\int_{T}t\wedge{\scriptstyle\bigstar}g=0.$ (89)
This gives $u=0$.
We have proved that the bilinear form (78) defines an injective map:
$A^{k}_{0}(T)\to\mathcal{P}_{\!-}^{r}\Lambda^{d-k}(T)^{\star}.$ (90)
(iv) We proceed to prove that the two spaces above have the same dimension.
For $k<d$, we have from [16, Equation (2.28)]:
$\dim G^{k}(T)=\dim\mathcal{P}_{\!-}^{r}\Lambda^{d-k}(T).$ (91)
Recall also the injectivity of the map $G^{k}(T)\to A^{k}_{0}(T)$, proved above. We get:
$\dim G^{k}(T)=\dim A^{k}_{0}(T)=\dim\mathcal{P}_{\!-}^{r}\Lambda^{d-k}(T).$ (92)
For $k=d$ we have more simply:
$\dim A^{d}(T)=\dim\mathcal{P}^{r}\Lambda^{d}(T)=\dim\mathcal{P}_{\!-}^{r}\Lambda^{0}(T).$ (93)
It follows that the bilinear form (78) is invertible.
(v) Now it follows from Proposition 3.5 that the FES admits extensions in the sense of [21, Definition 2.3] (or [20, Definition page 75]). The unisolvence of the degrees of freedom follows then from a general induction argument given for FES in [21, Proposition 2.5] (or [20, Proposition 5.35]). ∎
Proposition 3.10 (Computation of moments).
For elements $\in A^{k}(T)$ its moments against $\mathcal{P}^{r}\Lambda^{d-k}(T)$ can be computed inductively from the degrees of freedom.
Proof.
(i) Let $u\in A^{k}(T)$. Suppose that $w\in\mathcal{P}^{r}\Lambda^{d-k}(T)$ and that we want to compute:
$\int_{T}u\wedge w.$ (94)
We write as before:
$(d-k+r)w=\mathrm{d}\kappa_{T}w+\kappa_{T}\mathrm{d}w-Q_{r-1}w.$ (95)
On the right side the two last terms are degrees of freedom, so we concentrate on the first. We have:
$\int_{T}u\wedge\mathrm{d}\kappa_{T}w=\pm\int_{\partial T}u\wedge\kappa_{T}w\pm\int_{T}\mathrm{d}u\wedge\kappa_{T}w.$ (96)
(ii) As an intermediate step, we now treat the case where $\mathrm{d}u=0$. Then there is only the first term on the right side. It reduces to terms of the following form, where $T^{\prime}$ is a codimension one face of $T$:
$\int_{T^{\prime}}u\wedge(\kappa_{T}-\kappa_{T^{\prime}})w+\int_{T^{\prime}}u\wedge(\kappa_{T^{\prime}}w-P_{r}\kappa_{T^{\prime}}w)+\int_{T^{\prime}}u\wedge P_{r}\kappa_{T^{\prime}}w.$ (97)
Here $P_{r}$ is $L^{2}$ projection onto polynomial differential forms of polynomial degree $r$. Since $P_{r}$ is applied to polynomial forms (or degree at most $r+1$) on a polytope, it is computable.
In (97) the first and last terms are moments against $\mathcal{P}^{r}\Lambda^{d-1-k}(T)$, so amenable to an induction step. The middle term is zero, since it is a moment against $U^{d-1}(T^{\prime})$ representing a constraint.
This settles the intermediate step: we can compute the moments of $u$ against polynomials of degree $r$ in the case $\mathrm{d}u=0$, from the dofs of $u$, by induction on dimension.
(iii) In the general case we remark that $\mathrm{d}\mathrm{d}u=0$. Applying the intermediate step to $\mathrm{d}u$, its moments against polynomials of degree $r$ can be computed. Going back to (96) and looking at the last term, we write:
$\int_{T}\mathrm{d}u\wedge\kappa_{T}w=\int_{T}\mathrm{d}u\wedge(\kappa_{T}w-P_{r}\kappa_{T}w)+\int_{T}\mathrm{d}u\wedge P_{r}\kappa_{T}w.$ (98)
On the right side the first term is zero by construction, since $\kappa_{T}w-P_{r}\kappa_{T}w\in U^{d-k-1}$. The last term is computed by the above intermediate step.
In (96) we are left with the first term on the right side, for which (97) still applies. Again, this provides the beginnings of an induction procedure. ∎
From this proposition one can define a discrete computable $\mathrm{L}^{2}$ product on $A^{k}(T)$ as in (37), which is exact as in (38), when one of the fields is polynomial of degree $r$.